# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v4

@@ Abstract
Index Terms:
Federated Learning, Differential Privacy, Fairness
I Introduction
@@
Figure 1: Client drift caused by Non-IID.
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary [25]. Current fairness-aware machine learning algorithms, such as [10, 8], have not fully addressed the issue of client performance disparities in federated learning. As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity [28, 26, 50].
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data [53] or member inference attacks [42]. Even if only models or gradients are uploaded instead of original data, data leakage can still occur [35]. To enhance data privacy in FL, differential privacy (DP) [11] has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [1, 5, 12].
@@
    In addition, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
@@
    Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness. What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
II Related Work
II-A Differentially Private Federated Learning
Differentially Private Federated Learning (DPFL) has become widely used to enhance privacy protection in federated learning. Enayat et al. [45] dynamically adjusted compression rates in response to training loss changes to minimize communication rounds. Zhou et al. [52] examined the impact of local iterations when using the Laplace mechanism for perturbation. Fu et al. [13] introduced adaptive noise addition, customizing it based on loss fluctuations across clients and communication rounds. Ling et al. [31] balanced communication resources and privacy budgets by identifying the locally optimal number of iterations through convergence analysis. Raouf et al. [20] improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
II-B Fairness in DPFL
The above work in DPFL mainly focuses on model performance and privacy budget efficiency, without addressing fairness issues. In machine learning, a range of fairness concepts has been proposed, each emphasizing different aspects to meet diverse stakeholder needs. For instance, Good-Intent Fairness [38] and Selection Fairness [51] have been proposed. For a deeper understanding of these concepts, please refer to the relevant literature [7, 32, 34, 41].
In contrast, some recent studies have begun exploring group fairness within DPFL, aiming to reduce biases related to sensitive attributes such as gender and race [15, 39, 16, 17, 40]. Research suggests that adjusting parameters like clipping thresholds and noise multipliers can enhance group fairness in DPFL algorithms [16]. Under the framework of DP, adjacent datasets are defined as datasets $D$ and $D^{\prime}$ that differ by a single element, ensuring that an adversary cannot determine whether the output originates from $D$ or $D^{\prime}$. This mechanism significantly enhances the protection of attribute privacy and effectively reduces potential discrimination risks. Jagielski et al. [18] pioneered research on fair learning under DP constraints, addressing the need for predictive models to avoid discrimination based on specific attributes. They designed two learning algorithms that ensure both DP and equal opportunity, addressing challenges related to the collection of sensitive attributes. Tran et al. [44] utilized DP and Lagrangian duality to design neural networks capable of addressing fairness constraints while protecting the privacy of sensitive attributes. This approach ensures compliance with legal and ethical neutrality requirements when handling data related to gender, race, or age. Similarly, Ding et al. [9] proposed a classification model that combines a functional mechanism with decision boundary fairness to achieve both fairness and DP. This approach addresses privacy concerns related to training data and mitigates unfair practices stemming from attributes such as gender and race.
II-C Balanced Performance Fairness
Unlike the studies mentioned above, our research shifts focus to a relatively unexplored area within DPFL—balanced performance fairness. While group fairness emphasizes reducing performance biases related to sensitive attributes across different data types and is primarily used in centralized machine learning, balanced performance fairness, introduced by Li et al. [28], is more suited to federated learning. It emphasizes balancing the performance of global model across all clients, which is especially meaningful for data providers like companies and institutions. Building on the work of [28], Zhao et al. [50] proposed a strategy that replaces loss amplification with direct weight reallocation. Meanwhile, Li et al. [26] concentrated on improving fairness in personalized federated learning for local models.
Incorporating DP into model updates, with parameter adjustments such as clipping bounds and noise multipliers, can influence group fairness to some extent [16]. However, achieving balanced performance fairness necessitates adjustments to the direction of model updates. Therefore, determining how to update the model under DP constraints to ensure balanced performance fairness, while accounting for the effects of clipping and noise addition, presents a significant challenge. To the best of our knowledge, this is the first study to address balanced performance fairness within the context of DPFL.
III Preliminaries
In the beginning, Table I presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
TABLE I: Summary of main notations
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
III-A Federated Learning and Fairness
@@
$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
@@
(Balanced performance fairness [28]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
In this study, we employ the loss of model denote the performance. Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
III-B Differential Privacy
Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms. It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
@@ III-B Differential Privacy
(Differential Privacy [11]). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
@@
By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2. The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow. We use $l_{2}$ sensitivity as in the previous work.
@@
  • Firstly, computing the gradient for data sample $\xi_{j}$: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
@@
    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$ (4)
@@
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
@@
IV FedFair: Equipping FL with Fairness
@@
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
@@ 2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
@@ 8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
@@
15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
16       for batch $b\in\mathcal{B}_{i}$ do
@@ 21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in Section VI-B. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (7), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (7) as follows:
@@
The stochastic gradient descent by using Equation (IV) as show:
@@
  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
V FedFDP: Further Equipping FedFair with Differential Privacy
In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy. The flow of FedFDP algorithm is shown in Figure 2. It adds two additional processes to FedFair:
  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
Next, we discuss the details of these two strategies.
V-A Fair-clipping Strategy for Gradient
@@
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (10)
@@
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\tilde{g}^{i}_{t}$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
@@
V-B Adaptive Clipping Method for Loss
Since calculating $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, differential privacy must be incorporated to ensure the algorithm preserves privacy. As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $[0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, determining an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ prevents distortion of the original values but results in larger injected noise.

### v6

@@ Abstract
1 Introduction
@@
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary [22]. Current fairness-aware machine learning algorithms, such as [10, 8], have not fully addressed the issue of client performance disparities in federated learning. Due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity [25, 23, 42].
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data [43] or member inference attacks [35]. Even if only models or gradients are uploaded instead of original data, data leakage can still occur [31]. To enhance data privacy in FL, differential privacy (DP) [11] has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [12].
@@
    Furthermore, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
@@
    Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness. What’s more, we investigated the impact of clipping bound and noise multiplier on fairness.
In Section 2, we formalize our problem and introduce the three objectives of our method. Section 3 presents a fair federated learning approach—FedFair. In Section 4, we enhance FedFair by incorporating the DP guarantee, thereby arriving at the proposed FedFDP. A fairness parameter $\lambda$ is embedded within FedFDP, and in Section 5, we derive the analytical solution for the optimal $\lambda^{*}$ through convergence analysis and differential methods. The privacy analysis of FedFDP are examined in detail in Section 6. In Section 7, we conduct extensive experiments on three datasets, comparing six baseline algorithms to rigorously validate the effectiveness of both FedFair and FedFDP. Following the review of related work in Section 0.B, Section 9 provides a summary of our contributions. Due to space limitations, the detailed procedures of the convergence analysis are included in Appendix 0.A.
2 Problem Formulation
In this work, we propose FedFDP which aims to serve threefold goals as follows. Then, we will elaborate on and formalize how fairness and DP are defined in FL.
  • Goal 1 (Fairness): Strive to achieve federated learning with the highest possible fairness, where lower values of the federated fairness metric (Equation 2) indicate better fairness.
  • Goal 2 (Privacy): Implement federated learning under formal differential privacy guarantees (Equation 3).
  • Goal 3 (Utility): Ensure that the federated learning model retains strong predictive performance, particularly in terms of classification accuracy.
2.1 Fairness in FL
@@
$\mathbf{w}^{*}$ $=\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{\sum_{j}|D_{j}|}$, $D_{i}$ represents the dataset of client $i$.
@@
(Balanced performance fairness [25]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the NN devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the NN devices.
Differing from [28] that uses the variance of accuracy for "performance", we employ the variance of loss. This enables the metric to be incorporated directly into the objective function, facilitating more straightforward iterative optimization. Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
$\Psi(\mathbf{w})=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
2.2 DP guarantee in FL
DP is a formal mathematical model that quantifies privacy leakage in data analysis algorithms. It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
@@ 2.2 DP guarantee in FL
(DP [11]). $(\epsilon,\delta)$-DP is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single data sample, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
@@
By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2. The $l_{2}$-sensitivity determines how much noise is needed and is defined as follow.
@@
  • Firstly, computing the gradient for data sample $\xi_{j}$, donated by $\mathbf{g}^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
@@
    $\mathbf{\hat{g}}^{i,j}_{t}=\mathbf{g}^{i,j}_{t}/\max\left(1,\frac{\|\mathbf{g}^{i,j}_{t}\|}{C}\right),$ (4)
@@
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\cdot\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier, $\mathbf{I}$ is the identity matrix:
    $\mathbf{\tilde{g}}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\mathbf{\hat{g}}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
@@
3 FedFair: Equipping FL with Fairness
@@
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$;
@@ 2 SERVER
3 for $t=0,1,\cdots,T-1$ do
4    for $i=1,2,\cdots,N$ parallel do
@@ 8    $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
13 $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
@@
15 $\{\mathcal{B}_{i}\}\leftarrow$ Split $D_{i}$ to batches;
16 for batch $b\in\{\mathcal{B}_{i}\}$ do
@@ 21 $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
where $F(\mathbf{w}^{i}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round t1t-1, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focuses on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in Section 5.1. In addition, as shown in Figure 2, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (7), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (7) as follows:
@@
The stochastic gradient descent by using Equation (3) as show:
@@
  2. 2.
    Local (lines 1-1): For each batch $b$, we compute the dynamic learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
4 FedFDP: Further Equipping FedFair with Differential Privacy
In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy. The flow of FedFDP algorithm is shown in Figure 1. In addition to adding Gaussian noise to the gradients, we introduce two additional processes to balance fairness and DP: 1) A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness. 2) An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility. Next, we discuss the details of these two strategies.
4.1 Fair-clipping Strategy for Gradient
@@
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (10)
@@
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\mathbf{\tilde{g}}^{i}_{t}$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right]\,$, (12)
@@
4.2 Adaptive Clipping Method for Loss
Since calculating $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, differential privacy must be incorporated to ensure the algorithm preserves privacy. As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, determining an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. In federated learning, due to the heterogeneity of data held by different clients, $F_{i}(\mathbf{w}_{t}^{i})$ can vary significantly. A clipping norm that is too large introduces excessive noise, while one that is too small leads to aggressive clipping of gradient directions, impairing model performance. Therefore, an adaptive clipping strategy is necessary.